A web music store offers two versions of a popular song. The size of the standard version is 2.5 megabytes (MB). The size of the high-quality version is 4.9 MB. Yesterday, there were 710 downloads of the song, for a total download size of 2279 MB. How many downloads of the high-quality version were these?
Just trying to see what formula to Use...
Do I Divide 49. by 2,279?

Answer by Boreal(15235)

This one I did yesterday.
x= the standard version
y=the 4.9 version
x+y=710
2.5x + 4.9 y=2279
multiply first by (-2.5)
-2.5x-2.5y = -1775;; Now add the first and the second to eliminate x.
2.4 y=504;;; divide both sides by 2.4
y=210 of the 4.9 MB
x=500 of the 2.5 MB
Check
210*4.9=1029
500*2.5=1250
Total is 2279 , all units are MB